OpenStudy (anonymous):

You deposit $5000 into an account that earns 5% compounded annually. A friend deposits $4750 into an account that earns 4.95% annual interest, compounded continuously. Will your friend's balance ever equal yours?

OpenStudy (anonymous):

I tried making the equations for this: yours:\[A=5000(1+.05)^t\] friends: \[q=4750*e ^{0.0495*t}\]

OpenStudy (anonymous):

and graphed them to see when they would intersect but found they didn't but my homework says that's wrong. Not sure where to go with this now

OpenStudy (anonymous):

@ganeshie8 are you able to help with this one?

ganeshie8 (ganeshie8):

see if \(e^{0.0495} \gt 1+0.05\) is true

ganeshie8 (ganeshie8):

if it is, then the `continuous compounding` curve grows at a greater rate and overtakes the `annual compounding` curve at some point for sure

OpenStudy (anonymous):

well unless my equations are wrong they ran pretty consistently alongside each other when I graphed it

ganeshie8 (ganeshie8):

graphing is a good idea, but consider the possibility that they might cross much far away... half way before infinity ?

ganeshie8 (ganeshie8):

if the `continuous` growth rate is even a hair greater than the `annual compounding` growth rate, then the `continuous` wins in the long run

OpenStudy (anonymous):

not with the curves they were at. they would have to curve back around towards each other which isn't a quality of exponential equations. so I did something wrong somewher

ganeshie8 (ganeshie8):

Did u check if below is true : \(e^{0.0495} \gt 1+0.05\) ?

OpenStudy (anonymous):

yeah, it's > by like o.ooo7

ganeshie8 (ganeshie8):

lol thats more than a hair

ganeshie8 (ganeshie8):

so ur friend's balance is growing at a slightly greater rate than your balance

ganeshie8 (ganeshie8):

so he will over take u after few years

ganeshie8 (ganeshie8):

you may not be living to see it though... he reaches it in approximately 72 years : http://www.wolframalpha.com/input/?i=solve+5000%281%2B.05%29%5Et+%3D+4750*e+%5E%7B0.0495*t%7D

OpenStudy (anonymous):

so is 72 years when our accounts will be equal? I'm confused

ganeshie8 (ganeshie8):

yup thats what wolfram says

ganeshie8 (ganeshie8):

but u dont need to work that

ganeshie8 (ganeshie8):

the question is asking u only whether he will ever reach ur balance or not

ganeshie8 (ganeshie8):

for that checking the growth rate is sufficient

OpenStudy (anonymous):

no the question is asking WHEN he will equal mine

ganeshie8 (ganeshie8):

you better read the question again :/

ganeshie8 (ganeshie8):

or i must have overlooked something, let me go thru again lol

OpenStudy (anonymous):

that's what the question says but the answer box says " ____years (if he never reaches your balance enter NEVER)"

OpenStudy (anonymous):

72 was right. i just have no idea how you got it

OpenStudy (anonymous):

can you just show me how you got 72?

ganeshie8 (ganeshie8):

ohhk.. cool :)

ganeshie8 (ganeshie8):

you wanto knw when the balances in both accounts equal, so simply equate both the equations and solve for \(t\)

ganeshie8 (ganeshie8):

you equation : \(A=5000(1+.05)^t \) your friend's equation : \(q=4750*e ^{0.0495*t}\)

ganeshie8 (ganeshie8):

set them equal : \(5000(1+.05)^t=4750*e ^{0.0495*t}\)

ganeshie8 (ganeshie8):

solve

OpenStudy (anonymous):

oh okay, that makes more sense, thank you :D

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

wait, one more question. How would you go about solving that? Would you take the log of both sides? your link to wolfram was kinda confusing on how to solve it. I have to show my work or I don't get credit

ganeshie8 (ganeshie8):

\(5000(1+.05)^t=4750*e ^{0.0495*t}\) divide 4750 both sides : \(\dfrac{5000}{4750}(1.05)^t=e ^{0.0495*t}\) divide (1.05)^t both sides : \(\dfrac{5000}{4750}= \dfrac{e ^{0.0495*t}}{(1.05)^t}\) which is same as : \(\dfrac{5000}{4750}= \left(\dfrac{e ^{0.0495}}{1.05}\right)^t\)

ganeshie8 (ganeshie8):

Now you can take log both sides and isolate \(t\)

OpenStudy (anonymous):

Okay, i got it now.

OpenStudy (anonymous):

Thank you :)

ganeshie8 (ganeshie8):

u wlc :)